# Parameter-Specific Bias Diagnostics in Random-Effects Panel Data Models

### v4

@@ 5. Conclusion
We have illustrated how the bias test of Karl & Zimmerman (2021) can be used to detect instances where the mixed model (1) produces biased estimators for fixed effects. Besides considering bias rather than consistency, this diagnostic is narrower than the Hausman test in the sense that it targets the component of bias captured by $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$, i.e., bias operating through systematic alignment of the random effects with the weighting vector $\hat{\boldsymbol{\nu_{k}}}$ when the fixed- and random-effects design matrices are not orthogonal (with respect to $\mathbf{R}^{-1}$). In the stochastic-$\mathbf{Z}$} setting emphasized by Karl & Zimmerman (2021), such alignment is induced by dependence between $\mathbf{Z}$} and $\boldsymbol{\eta}$. In contrast, the Hausman test is sensitive to any potential source of inconsistency in the random-effects estimator of the fixed effects, provided the fixed-effects estimator is consistent. Nevertheless, the bias diagnostic can be computed from the fitted random-effects model without changing the original RE specification and generates summaries for each fixed effect or contrast. These features make it a potential complement to the Hausman test when assessing potential misspecification in linear mixed models.

### v5

@@ 5. Conclusion
We have illustrated how the diagnostic of Karl & Zimmerman (2021) can be used to flag fixed effects and contrasts for which a fitted mixed model may exhibit non-negligible finite-sample bias. Relative to the Hausman test, the diagnostic is narrower: it targets the coefficient-specific alignment term $\mathbf{E}[\hat{\boldsymbol{\nu_{k}}}^{\prime}\boldsymbol{\eta}]$ rather than every possible source of inconsistency in the RE estimator. Its formal sampling interpretation is sharpest in the stochastic-$\mathbf{Z}$} settings emphasized by Karl & Zimmerman (2021), where nonrandom assignment can induce alignment between the random-effects design and the random effects themselves. In more conventional panel settings, we view the resulting statistic primarily as a descriptive complement to FE–RE comparisons. Even so, the diagnostic is attractive because it can be computed from a single fitted RE model and reported for any fixed effect or contrast of substantive interest.